What is the degree of the polynomial -4x(to the third power)+2x(to the second power)y-5xy(to the fourth power)
and also
which of the following is a monomial with a degree of 7
a.-2x(to the third power)y( to the second power)
b.4x(to the fourth power)+ 5xy( to the second power)
c.7x
d.x( to the fifth power) yz

Answer by mathchemprofessor(65)

5x.y^4=5th degree.
d)x^5.yz=7 (adding the exponents of x,y andz as 5+1+1=7)